OpenStudy (anonymous):

Consider the function:y = x^2+2x. What is the derivative dy/dx of this function

OpenStudy (anonymous):

2x + 2

OpenStudy (anonymous):

2x+2

OpenStudy (anonymous):

FIRST! haha the general rule is: \[f(x)=x^{a}\] \[f \prime(x)=ax ^{a-1}\]

OpenStudy (anonymous):

firstly: y=f(x) = x^2 +2x then using the rule bananamoes gave which is correct! It follows f ' (x) = 2x^(2-1) + 2x^(1-1) = 2x + 2 as x^0 = 1